### v1

# CReFT-CAD: Boosting Orthographic Projection Reasoning for CAD via Reinforcement Fine-Tuning

Abstract
Computer-Aided Design (CAD) plays a pivotal role in industrial manufacturing. Orthographic projection reasoning underpins the entire CAD workflow, encompassing design, manufacturing, and simulation. However, prevailing deep-learning approaches employ standard 3D reconstruction pipelines as an alternative, which often introduce imprecise dimensions and limit the parametric editability required for CAD workflows. Recently, some researchers adopt vision–language models (VLMs), particularly supervised fine-tuning (SFT), to tackle CAD-related challenges. SFT shows promise but often devolves into pattern memorization, yielding poor out-of-distribution performance on complex reasoning tasks. To address these gaps, we introduce CReFT-CAD, a two-stage fine-tuning paradigm that first employs a curriculum-driven reinforcement learning stage with difficulty-aware rewards to build reasoning ability steadily, and then applies supervised post-tuning to hone instruction following and semantic extraction. Complementing this, we release TriView2CAD, the first large-scale, open-source benchmark for orthographic projection reasoning, comprising 200,000 synthetic and 3,000 real-world orthographic projections with precise dimension annotations and six interoperable data modalities. We benchmark leading VLMs on orthographic projection reasoning and demonstrate that CReFT-CAD substantially improves reasoning accuracy and out-of-distribution generalizability in real-world scenarios, offering valuable insights for advancing CAD reasoning research. The code and adopted datasets are available at https://github.com/KeNiu042/CReFT-CAD.
footnotetext: Equal contribution. 🖂 Corresponding authors.
1 Introduction
Computer-Aided Design (CAD) is now integral to industrial product development, driving design, manufacturing, and simulation workflows. In the design phase, engineers use CAD drawings or rasterized orthographic projections for their precision and easy editability. During manufacturing, these drawings are converted into constraint-based parameter tables, and for simulation, they yield boundary-representation (B-Rep) data or textual geometry descriptions. A truly user-centric pipeline thus requires accurate semantic parsing of orthographic projection: automated extraction of parameter tables directly produces 3D models that meet stringent manufacturing and simulation standards. By contrast, existing reverse-engineering methods rely on expensive 2D/3D scanning hardware and labor-intensive post-processing, severely limiting scalability and broad industrial adoption.
With the advent of deep learning, orthographic projection reasoning has typically been cast as a standard 3D reconstruction pipeline: rasterized drawings are processed by models that output B-Rep structures [1, 2], point clouds [3, 4], or meshes [5]. However, two critical shortcomings remain. First, industrial drawings allow virtually zero tolerance for missing or erroneous components: pixel-level discretization errors in reconstruction can propagate into the 3D model, causing failures in downstream manufacturing and simulation. Second, practical CAD workflows demand parametric editability and precise semantic alignment—capabilities that common 3D representations lack. Although some methods generate CAD command sequences to restore editability, they still overlook the nuanced semantics of orthographic views and thus fail to map 2D drawing features to their intended 3D counterparts reliably.
Recent works leverage the modality alignment and semantic reasoning strengths of vision–language models (VLMs) [6, 7, 8, 9, 10, 11], applying supervised fine-tuning (SFT) to address CAD challenges[12, 13, 14]. However, orthographic projection reasoning demands genuine inferential ability, and task-specific SFT often leads to memorization and out-of-distribution (OOD) degradation rather than deeper reasoning [15]. Inspired by DeepSeek R1-Zero’s Group Relative Policy Optimization (GRPO) [16], we propose Curriculum-driven Reinforcement Fine-tuning for CAD (CReFT-CAD),
CReFT-CAD is a versatile framework that assists engineers in achieving orthographic projection reasoning throughout the entire lifecycle of industrial design, from design and manufacturing to simulation, in an interactive manner, as shown in Fig. 1. CReFT-CAD is a two-stage paradigm that combines reinforcement learning and supervised tuning. Curriculum-driven Reinforcement Fine-tuning: We introduce a difficulty-aware reward mechanism that incrementally escalates task complexity, gradually exposing the model to more complex reasoning challenges and promoting stable policy optimization. Supervised Post-tuning: Building on the reinforced model, we further refine its instruction-following and reasoning capabilities via multi-task SFT. Our qualitative and quantitative evaluations demonstrate that CReFT-CAD delivers robust orthographic projection reasoning and strong generalization in OOD scenarios. Another major obstacle in orthographic projection reasoning has been the lack of open-source datasets with high-fidelity annotations. To fill this gap, we introduce TriView2CAD, the first large-scale benchmark tailored for industrial CAD pipelines. Leveraging real-world design archives, we employ a constraint-driven synthesis pipeline to generate 200,000 synthetic and 3,000 real-world three-view sets, each annotated with precise, one-to-one dimension labels tied to their corresponding geometric primitives. This design enables models to extract quantitative measurements from rasterized drawings and enforce exact geometric constraints for simulation-ready 3D reconstruction.
Figure 1: (a) (b) show the results of parameterization tasks of orthographic projection reasoning using Qwen2.5-VL and Deepseek without tuning. (c)–(f) illustrate CReFT-CAD’s capabilities across various orthographic projection reasoning tasks.
2 Related Work
2.1 Orthographic projection reasoning for CAD
Understanding and reasoning with orthographic projection is a fundamental and longstanding challenge in CAD-related tasks [17, 18, 19]. With the development of 3D reconstruction methods, existing orthographic projection reasoning methods have largely focused on standard 3D reconstruction pipelines. SPARE3D [20] introduces a dataset designed to evaluate the spatial reasoning capabilities of AI systems through 2D line drawings of 3D objects. Contrastive-SPARE3D [21] propose a self-supervised binary classification network that helps to learn 3D object line drawings representations that are detail-sensitive and view-invariant. PlankAssembly [22] introduces a transformer-based sequence generation model that learns flexible mappings between inputs and outputs. IsoTGAN [23] introduces a novel Gaussian-enhanced Euclidean attention mechanism and a geometric constraint loss function to enhance local image features further. GaussianCAD [24] employs a custom sparse-view 3D reconstruction method, eliminating the reliance on vector CAD sketches and 3D real-world data.
2.2 Applications of LVLMs in CAD-Related tasks
Recent work on applying vision–language models (VLMs) to CAD tasks can be grouped into three main categories: 1) Direct Adaptation of Pretrained VLMs. CAD-Recode [25] leverages a VLM to translate raw point-cloud inputs into executable Python code. Both LEAM [26] and LLM4CAD [27] generate CAD models by jointly leveraging textual descriptions and image inputs. CAD-Assistant [28] introduces a novel approach that leverages VLMs as planners, where the model interacts with Python APIs to accomplish various CAD tasks. 2) Supervised Fine-Tuning (SFT). CAD2Program [12] generates 3D parametric models from 2D CAD drawings. CAD-MLLM [29] enables parametric CAD modeling from text, images, and point clouds, and proposes the Omni-CAD dataset. 3) Reinforcement Learning–Based Tunings. RLCAD [30] proposes a reinforcement learning environment to generate complex CAD models. CADCrafter [31] presents an image-to-CAD model generation framework, fine-tuned via Direct Preference Optimization (DPO) to translate input images directly into executable CAD representations.
3 TriView2CAD: Benchmarking orthographic projection reasoning
As previously noted, a critical barrier in CAD research is the absence of an open-source dataset with high-fidelity annotations for orthographic projection reasoning. Existing benchmarks focus on 3D reconstruction from rasterized 2D views, yet diverge from real-world engineering drawing interpretation in three key aspects:
Lack of precise dimensional annotations. Existing orthographic projection benchmarks provide only rasterized drawings without precise dimension annotations. It causes models to prioritize relative scale over exact measurements, and pixel-level errors in raster images can easily propagate into the reconstructed 3D model, resulting in inaccuracies. Additionally, the limited editability of reconstructed 3D models precludes corrective adjustments during design and simulation.
Lack of explicit logical reasoning tasks. No current dataset evaluates a model’s inferential capabilities within or across orthographic views. Consequently, models trained on these benchmarks learn pixel-level reconstruction but fail to handle real-world challenges such as inferring omitted annotations or exploiting structural symmetries.
Lack of essential CAD modalities. Most datasets include only one or two modalities—typically raster images and, at best, a single 3D format (meshes or point clouds). This limits coverage omits critical data representations (e.g., parameter tables, vector CAD files, executable commands, STEP/B-Rep) needed for end-to-end CAD design, manufacturing, and simulation pipelines.
3.1 Dataset construction
In this work, we focus on prefabricated bridge piers due to their inherently modular composition of repeated, standardized components, which yields a compact yet expressive parameter space for dataset synthesis. Moreover, prefabricated piers are ubiquitous in modern infrastructure and supported by extensive archives of CAD drawings and manufacturing documentation. TriView2CAD comprises 200,000 synthetic samples and 3,000 real-world samples. The synthetic data are split 80/20 into training and test sets, while, due to the scarcity of real-world engineering drawings, all 3,000 real-world samples are reserved exclusively for testing. The synthetic test set is used to evaluate in-domain performance. As shown in Fig. 3, real-world images typically exhibit more complex structures, primarily due to redundant annotation lines causing occlusions, overlap with other components, and the inclusion of non-numeric annotations, serving to assess out-of-distribution generalization in authentic CAD scenarios. To mirror end-to-end industrial workflows, TriView2CAD provides six interconnected data modalities per sample—structured parameter tables (JSON), vector CAD drawings (DXF), raster images (PNG), executable modeling scripts, and two standard 3D formats (STEP and B-Rep). This comprehensive suite supports tasks from early design to downstream manufacturing and simulation. As shown in Fig. 5, we employ a constraint-driven synthesis pipeline. The detailed steps are as follows:
Figure 2: Constraint-driven synthesis pipeline for TriView2CAD.
Figure 3: Examples of Real-World Orthographic Projections and its 3D model.
Step 1: Constraint-Guided Parameter Sampling We first analyze real-world CAD drawings to define a 15-dimensional parameter space subject to two constraint classes. Intra-View Constraints enforce both topological closure and physical validity within each projection: every component must form a gap- and overlap-free contour, and paired dimensions must satisfy domain-specific engineering constraints (e.g., “Cross-Bridge Pier Spacing” < “Cap Beam Cross-Bridge Dimension”). Inter-View Constraints ensure cross-projection consistency by requiring every measurement annotated in one view to exactly match its counterpart (height, width, and depth) in the other views. By embedding these constraints into our sampling pipeline, we guarantee that every synthetic design is geometrically coherent and adheres to real-world engineering practices.
Step 2: 2D Drawing and Raster Image Generation We employ the ezdxf library to translate each sampled parameter vector into a vectorized 2D CAD drawing (DXF), mapping numeric dimensions to primitives such as lines, circles, and arcs to form coherent orthographic projections. We leverage ezdxf’s layers-separating function to partition primitives into semantic layers, thereby enhancing the editability of the resulting CAD drawings. Each DXF file is then loaded into FreeCAD [32], where we capture high-resolution screenshots to produce three orthographic views (front, top, and side).
Step 3: 3D Model Generation We employ the FreeCAD Python API to reconstruct 3D model from its parameter vector. Mirroring our 2D pipeline, we decompose the model into individual entities and generate scripted commands that accurately position each primitive according to the intra- and inter-view constraints established in Step 1. By executing these scripts, we automatically produce industry-standard STEP and B-Rep files, closing the loop from parameter sampling to editable, simulation-ready CAD geometry.
3.2 Evaluation of TriView2CAD
We benchmark seven leading vision–language models on TriView2CAD to perform high-precision orthographic projection reasoning. The evaluation results are presented in Sec. 3.2. Reflecting real-world CAD workflows, we define three complementary evaluation tasks to rigorously test a model’s ability to extract precise dimension annotations from rasterized orthographic projection and infer the underlying geometric relationships. (i) Dimension recognition and pairing Models identify each annotated dimension in a single orthographic view and map it to its corresponding geometric feature. (ii) Primitive counting Models count instances of specified CAD elements (e.g., number of pier columns), testing their ability to parse structural composition. (iii) Composite Parameter Computation Models compute engineering-critical derived quantities—for example, the Cross-Bridge Pier Spacing is defined as the sum of the Pier Column Cross-Bridge Dimension and the Pile Spacing. Overall, we define a 15-dimensional parameter space comprising 6 recognition parameters, 3 counting parameters, and 6 composite calculation parameters. During evaluation, we treat each individual parameter as a separate prediction. The overall accuracy is then computed as the total number of correctly predicted parameters divided by the total number of parameters across all test samples. This evaluation suite exposes each model’s strengths and weaknesses in handling the tightly coupled semantic and quantitative demands of orthographic projection CAD reasoning.
4 Methodology
Inspired by the success of reinforcement learning–based tuning methods [33, 34], we present Progressive Reinforcement Fine-tuning for CAD (CReFT-CAD), a orthographic-projection reasoning framework. CReFT-CAD combines a ViT-based visual encoder with Qwen2.5 [35]. First, Progressive Reinforcement Fine-tuning employs a curriculum of three progressively harder reasoning tasks under a difficulty-aware reward scheme, enabling the model to move beyond rote pattern matching toward robust reasoning while improving OOD resilience. Second, Supervised Post-tuning refines the model’s instruction-following to support interactive, real-world CAD queries.
4.1 Curriculum-driven reinforcement fine-tuning
Orthographic projection reasoning is a complex reasoning task that encompasses not only dimension recognition and pairing but also primitive counting and more intricate composite parameter computation tasks. Building on GRPO, we propose a Curriculum-driven Reinforcement Fine-tuning (CReFT) strategy, which incrementally exposes the model to progressively more complex training tasks to enhance its reasoning and generalization capabilities. Our tasks design mirrors the core steps of expert engineers in verifying CAD drawings, incorporating dichotomous choice task, multiple choice task, and parameterization tasks based on chain-of-thought (CoT). For each task, we design a tailored, efficient reward function to ensure the model can steadily and stably optimize its reasoning performance.
Figure 4: Diagram design of three tasks in curriculum-driven reinforcement fine-tuning.
Data engine. We leverage 160,000 image–text instruction pairs using the synthetic training set of TriView2CAD, where each prompt includes a given orthographic projection and all 15 parameter key-value pairs, as shown in Fig. 4. To ensure a balanced training signal across all three tasks, 50% of the training responses are entirely correct, while the remaining 50% contain errors of varying magnitude. Specifically, for the dichotomous choice task, we ensure that each negative response contains $n$ erroneous parameter values, where $n$ follows a uniform distribution $N\sim\text{Uniform}(a,b)$, where $a=1$ and $b=15$.
For the multiple choice task, $p$ parameter values are masked in each image-text instruction pair. The masking mechanism ensures the diversity of options. This approach prevents the model from simply deriving the answer through parameterization, bypassing the complex multiple choice reasoning required in the task. Instead, it forces the model to engage in decision-making actively, considering multiple potential solutions and selecting the correct responses based on a deeper understanding. For each instruction pair, unmasked values are correct, while for the incorrect parameter value lists, all unmasked values, except for randomly selected $q$ erroneous lists. Both the masked values $p$ and the number of incorrect values $q$ follow a normal distribution, introducing a smooth and gradual variation in the difficulty of the learning task. This variation helps the model to learn in a more structured and progressive manner.
For the parameterization task, we construct image-text instruction pairs based on Chain-of-Thought (CoT) reasoning. Specifically, as for Composite Parameter Computation, expert knowledge guides the model through a step-by-step reasoning guidance to arrive at the correct output. The step-by-step workflow of reasoning guidance is as follow: 1) recognize parameters required for the computation of the composite parameter; 2) provide the calculation formula for the composite parameter; 3) generate the predicted result based on the formula and the factor parameters
Training strategy. Given that our task involves highly customized inputs and outputs rather than a general-purpose problem, directly applying a pretrained VLM for reasoning tasks and generating required outputs proves ineffective. As a result, the model cannot effectively execute the reinforcement learning reward mechanism, as the reward will always be 0. To address this, prior to executing each training task, we pre-sample a subset of image-text instruction pairs to "warm up" the model, facilitating the initialization of the reinforcement learning (RL) reward mechanism.
4.1.1 Task 1: Dichotomous choice task
As for this task, the model’s output is constrained to dichotomous responses, either "yes" or "no." A "yes" answer is given only when the entire 15-dimensional parameter space is completely correct, while any error, even in a single parameter, results in a "no" response. This strict criterion ensures that the model is incentivized to produce fully accurate parameterizations. To align with this setup, the reward function is defined as follows:
$R_{\text{P1}}=\begin{cases}1,&\text{ if all parameters are correct}\\ 0,&\text{otherwise}\end{cases}$ (1)
4.1.2 Task 2: Multiple choice task
Training strategy. In the task 2, we introduce multiple choice tasks to further increase the complexity of decision-making. These tasks require the model to select from multiple parameter value lists, which significantly enhancing the cognitive load compared to binary decisions. Building on the first stage, we pre-sample a subset of image-text instruction pairs to perform instruction-following training on the LLM. This training allows the model to generate structured outputs in the form of multiple choice questions, enabling the reinforcement learning reward mechanism. Additionally, we designed a stringent reward function that encourages the model to maximize the utilization of image-text instruction pairs, ensuring that it extracts and applies all relevant information from both the visual and textual components. The reward function incentivizes the model to select the most accurate option from the provided set of choices. Any incorrect selection results in a reward of 0, thereby pushing the model to focus on precision and reliability in its decision-making process.
Let $S_{\text{correct}}$ be the set of correct parameter value lists, $S_{\text{incorrect}}$ be the set of incorrect parameter value lists, and $S_{\text{Selected}}$ be the set of selected parameter value lists by the model. The reward function $R_{p2}$ is defined as:
$R_{\text{P2}}=\begin{cases}1&\text{if }S_{\text{selected}}=S_{\text{correct}}\\ 0.2&\text{if }S_{\text{selected}}\supseteq S_{\text{correct}}\text{ but }S_{\text{selected}}\neq S_{\text{correct}}\\ 0&\text{if }(S_{\text{selected}}\cap S_{\text{correct}}\neq\emptyset)\text{ and }(S_{\text{selected}}\cap S_{\text{incorrect}}\neq\emptyset)\\ 0&\text{if }S_{\text{selected}}=S_{\text{incorrect}}\end{cases}$ (2)
Where $S_{\text{Selected}}=S_{\text{correct}}$ means the model selects only the correct parameter value lists; $S_{\text{Selected}}\supseteq S_{\text{correct}}\text{ but }S_{\text{Selected}}\neq S_{\text{correct}}$ means the model selects all correct lists but fails to select all of them; $S_{\text{Selected}}\cap S_{\text{correct}}\neq\emptyset$ and $S_{\text{Selected}}\cap S_{\text{incorrect}}\neq\emptyset$ means the model selects both correct and incorrect parameter value lists; $S_{\text{Selected}}=S_{\text{incorrect}}$ means the model selects incorrect parameter value lists.
4.1.3 Task 3: Parameterization tasks based on chain-of-thought
Training strategy. In the training process, we begin with the model trained in task 2, and sample synthetic data to construct standard orthographic projection reasoning question-answer pairs. These pairs are then used to perform small-batch fine-tuning. Once the model is fine-tuned, we assess the difficulty of predicting the 15-dimensional parameters by evaluating the model’s performance on a test set. The prediction accuracy for each parameter is categorized into three difficulty levels: Easy Attributes: Attributes with an accuracy greater than 0.8. Medium Attributes: Attributes with accuracy between 0.2 and 0.8. Difficult Attributes: Attributes with accuracy less than 0.2. This difficulty classification is then used to design the reward function The reward for each correctly answered attribute is determined by its difficulty level as follows:
$R_{\text{P3}}=\begin{cases}1&\text{if the model correctly predicts an easy attribute }\\ 1.5&\text{if the model correctly predicts a medium attribute }\\ 2&\text{if the model correctly predicts a hard attribute }\\ 0&\text{if the model's prediction is incorrect }\end{cases}$ (3)
This reward function ensures that the model is motivated to gradually improve across all difficulty levels, with greater emphasis on mastering the more challenging tasks, thus enhancing the model’s overall reasoning capabilities in orthographic projection tasks.
4.2 Supervised post-tuning
After the CReFT stage, where the model developed robust reasoning and generalization capabilities, the focus of the Supervised Post-tuning task shifts to enhancing the model’s instruction-following ability. During this stage, we rethink the industrial design application context and, with a user-centric approach, reformulate tasks in the style of Visual Question Answering (VQA). By leveraging multimodal language models, we design specific tasks that guide the model to output a comprehensive list of parameters and make critical judgments based on orthographic projection. In this stage, the model is trained to process the following tasks: 1) Full Parameter List Output Task: The goal is to generate a fully accurate set of parameter key-value pairs, which is the most common format for outputs during the simulation phase. 2) Parameter Validation Task: The task is to compare the extracted parameter key-value pairs with the accurate parameter key-value pairs in the orthographic projection to determine whether the extracted parameters are correct. 3) View Matching Task: The model is given two individual orthographic projection views. The model must, based on parameter constraints and rasterized images, assess whether the views correspond to the same overall 3D object or design. 4) Component Counting Task: The model is required to identify and count specific components in the drawing, such as bridge piers or columns, based on the given parameter constraints. The output of this task helps to quantify the number of elements in the design, aiding in material estimation and component verification.
By employing this post-tuning phase, we enhance the model’s ability to interpret and respond to complex design queries, thereby improving its interactive reasoning abilities and making it more applicable to real-world industrial design workflows.
5 Experiments
Implementation Details. All experiments are conducted on NVIDIA A100 GPUs. Most experiments use GOT-OCR2.0 as the base model, trained on a single server equipped with 8 A100 GPUs and a batch size of 64. We use the AdamW optimizer with a cosine annealing scheduler. The hyperparameters are configured as follows: (1) Learning rate: 1e-6 for RL (GRPO) training and 2e-5 for baseline SFT experiments. (2) Maximum input image size: 2,483,776 pixels. (3) Total GRPO training steps: 1500.
Datasets and Metrics. We evaluate both the leading vision–language models (VLMs) and the proposed method on TriView2CAD, following the evaluation strategy outlined in Sec. 3.2. The synthetic test set is used to assess in-domain accuracy. Due to the difficulty in obtaining real-world data, all 3,000 real-world samples are reserved exclusively for testing, allowing us to evaluate out-of-distribution (OOD) generalization in real-world scenarios.
We design four distinct prompt configurations to probe their orthographic projection reasoning capabilities: Test image-Only: Models receive only the target rasterized drawings. Test image + Reference Image: In addition to the target drawings, models are given a reference image—rendered on the same geometry layer—with all dimension names to supply semantic alignment. Test image + Answered Pair: Inputs consist of the target drawings plus a correctly paired raster image and its parameter vector, providing an exemplar mapping between visual features and quantitative parameters. Test image + Atttribute Explanation: Models are supplied with the target drawings alongside a detailed textual interpretation authored by professional engineers.
Given that orthographic projection reasoning involves complex multi-step inference and current pretrained VLMs limit exposure to CAD drawings, we incorporate explicit reasoning guidance into each prompt format to encourage clear reasoning trajectories. For each of the four different prompt input formats, we outline detailed step-by-step instructions for completing the task. These steps are designed to guide the model through the reasoning process in a structured manner. The complete set of instructions for each format is provided in the supplementary materials.
Table 1: Performance comparison of various VLMs on orthographic projection reasoning tasks.
Prompts Without Reasoning Guidance Reasoning Guidance
Test Img +Reference +Answered +Attribute Test Img +Reference +Answered +Attribute
Img Pair Explanation Img Pair Explanation
Phi-3.5-Vision [36] 4.22 8.95 8.05 6.26 11.32 14.40 16.16 15.79
LLaVA-OneVision [37] 9.16 16.06 15.65 14.84 9.10 16.81 16.21 20.53
DeepSeek-VL [38] 8.16 22.68 20.03 12.65 14.02 20.31 24.62 25.45
InternVL2.5 [39] 15.79 18.82 23.30 17.16 15.63 22.35 24.47 23.73
InternVL3 [39] 15.46 17.90 22.44 17.91 15.81 21.98 23.58 17.05
Qwen2.5-Omni [40] 23.43 28.89 28.74 26.50 26.12 30.93 29.40 35.71
Qwen2.5-VL [41] 24.54 30.76 30.47 25.86 24.54 32.78 33.64 38.88
Ours 80.86 82.99 83.24 82.67 81.35 83.11 82.87 84.03
5.1 Performance on in-domain test set.
The experimental results, as shown in Tab. 3, demonstrate that our method significantly outperforms existing VLMs on orthographic projection reasoning tasks across all prompt formats. Specifically, in the Without Reasoning Guidance setting, the "+Answered Pair" prompt achieves the highest accuracy of 83.24%. In the Reasoning Guidance setting, our model achieves an accuracy of 84.03% in the "+Attribute Explanation" prompt. Furthermore, our method exhibits strong performance even in the "Test Image Only" condition, where we achieve 80.86%. This improvement highlights the effectiveness of our approach.
It is worth noting that the performance improvement from "without reasoning guidance" to "with reasoning guidance" is not as significant. This is primarily because, through our CReFT approach, the model learns effective reasoning strategies via the difficulty-aware reward mechanism, thereby reducing the need for additional guidance to achieve further significant gains. Based on our benchmarking of the seven leading VLMs, we present several important findings, as follows: 1) This reasoning task remains a tough challenge for pretrained VLMs. Overall, orthographic projection reasoning entails not only reading textual dimensions and visual features, but also matching annotations to geometry primitives, counting structural elements, and computing composite parameters. 2) Different prompt formats influence the performance. In both the Without Reasoning Guidance and With Reasoning Guidance settings, different prompt formats lead to varying degrees of performance improvement. 3)Introducing reasoning guidance yields consistent gains across all seven VLMs. Introducing appropriate reasoning guidance significantly enhances the model’s ability to perform reasoning tasks effectively. Further analysis regarding the three findings are presented in the supplementary materials.
5.2 Performance on real-world test set.
We utilize all 3,000 real-world samples to evaluate out-of-distribution (OOD) generalization in real-world scenarios. The baseline model, Qwen2.5-VL, achieves an accuracy of 13.47%, which is lower than its performance on synthetic data (24.54%). This performance drop is primarily due to the more complex structure and occlusion relationships present in the real-world data, as discussed in Sec. 3.1, which increases the difficulty of the task. We also compare the results with a model trained using a SFT strategy, which achieve an accuracy of 36.15%. This improvement can be attributed to significant gains in Primitive Counting tasks. However, the model still struggle with Composite Parameter Computation, highlighting the limitations of traditional SFT. Our method, using CReFT, achieves an accuracy of 46.67%, demonstrating that our method can effectively address OOD generalization and lead to significant improvements on more complex real-world data.
Table 2: Ablation study results.
Method w/o CoT with CoT
Task3 46.16 74.15
Task1 + Task3 46.05 77.90
Task2 + Task3 30.71 67.75
Task1 + Task2 + Task3 46.24 81.35
5.3 Abalation study
With CoT VS without CoT. The results from Tab. 2 reveal a significant improvement in performance when incorporating chain-of-thought (CoT) reasoning. Specifically, the accuracy for individual Task 3 increase from 46.16% to 74.15% after adding CoT , demonstrating a clear performance boost. When combining multiple training tasks, such as Task 1 + Task 3 and Task 2 + Task 3, the improvements are even more pronounced, with accuracies reaching 77.90% and 67.75%, respectively. The most notable enhancement is observed in the combination of Task 1 + Task 2 + Task 3, where the accuracy increases from 46.24% to 81.35%. The primary source of these improvements lies in how CoT helps break down complex Composite Parameter Computation tasks into more manageable, incremental steps. The CoT mechanism, however, guides the model through each individual component of the computation, allowing for a step-by-step approach that improves the model’s reasoning capability.
Three training tasks. Notably, in the CoT setting, combining Task 1 and Task 3 provides a performance boost, increasing accuracy from 74.15% to 77.90%. However, when Task 2 is added, the model’s performance decreases. This decline can be attributed to the design of Task 2, which includes a substantial number of masked elements within the instruction data. By exposing the model to higher uncertainty, Task 2 forces the model to contend with incomplete or ambiguous information, which ultimately hinders its ability to perform effectively on the task. On the other hand, when all three tasks (Task 1 + Task 2 + Task 3) are used together, the model’s reasoning ability is fully activated, resulting in a significant performance improvement. In contrast, in the without CoT setting, the model’s performance is constrained by its inability to effectively handle multi-step reasoning tasks, where even the best-performing task combination achieves a maximum accuracy of only 46.24%.
6 Conclusion and Limitation
In this work, we address challenges in orthographic projection reasoning for Computer-Aided Design (CAD) by proposing CReFT-CAD, an innovative two-stage fine-tuning paradigm. We demonstrated that the integration of curriculum-driven reinforcement learning (RL) with difficulty-aware rewards effectively enhances the model’s reasoning ability. To further advance the field, we introduce TriView2CAD, the first large-scale, open-source benchmark specifically designed for orthographic projection reasoning. With its comprehensive dataset of 200,000 synthetic and 3,000 real-world samples, each annotated with precise dimensions and accompanied by six interoperable data modalities. In conclusion, this work lays the groundwork for more robust, scalable solutions to CAD-related challenges, providing both a novel training paradigm and a rich dataset that can foster future research and innovation in the field.
Limitation. The accuracy on complex real-world scenarios has not yet reached a very high level. As such, the model’s generalization ability remains a key area for further development. In future work, we plan to explore techniques to enhance the model’s robustness and adaptability.
Appendix A The complete reasoning guidance for each input format
For the Test Image-Only format:
  • Identify all annotated numbers in the orthographic projection, including dimensions, spacing, and height.
  • Interpret their meanings based on position, orientation, and surrounding context.
  • Directly assign values to some parameters, count graphical elements for quantity parameters, and compute spacing parameters based on position.
For the +Reference Image format:
  • Use the reference template to understand the correspondence between primitive and parameters, and identify annotated numbers in the target image.
  • Assign parameters based on position and template primitive, counting graphical primitive for quantities.
  • Calculate parameters based on positional relationships or template definitions, and output the full set of parameter values.
For the +Answered Image format:
  • Learn the correspondence between structures and parameters from the example image.
  • Identify and interpret annotated numbers in the target image, considering their position and similarity to the example.
  • Assign values to parameters, count graphical primitive for quantities, calculate parameters, and apply necessary constraints to produce final results.
For the +Attribute Explanation format:
  • Clarify parameter definitions and their geometric meanings, including component types and their properties.
  • Identify annotated numbers in the orthographic projection and infer their corresponding parameters.
  • Assign values to parameters, compute quantities, derive spacing values from positional relationships, and output the complete parameter set with necessary constraints.
Appendix B Further analysis regarding the three findings
Table 3: Performance comparison of various VLMs on orthographic projection reasoning tasks.
Prompts Without Reasoning Guidance Reasoning Guidance
Test Img +Reference +Answered +Attribute Test Img +Reference +Answered +Attribute
Img Pair Explanation Img Pair Explanation
Phi-3.5-Vision 4.22 8.95 8.05 6.26 11.32 14.40 16.16 15.79
LLaVA-OneVision 9.16 16.06 15.65 14.84 9.10 16.81 16.21 20.53
DeepSeek-VL 8.16 22.68 20.03 12.65 14.02 20.31 24.62 25.45
InternVL2.5 15.79 18.82 23.30 17.16 15.63 22.35 24.47 23.73
InternVL3 15.46 17.90 22.44 17.91 15.81 21.98 23.58 17.05
Qwen2.5-Omni 23.43 28.89 28.74 26.50 26.12 30.93 29.40 35.71
Qwen2.5-VL 24.54 30.76 30.47 25.86 24.54 32.78 33.64 38.88
Ours 80.86 82.99 83.24 82.67 81.35 83.11 82.87 84.03
1) Remains a Tough Challenge for pretrained VLMs. Tab. 3 reports performance across four prompt formats without and with reasoning guidance. Overall, orthographic projection reasoning entails not only reading textual dimensions and visual features, but also matching annotations to geometry primitives, counting structural elements, and computing composite parameters. Consequently, none of the off-the-shelf models fully masters this composite task. Qwen2.5-VL [42] achieves the highest accuracy—38.88% under reasoning guidance with the Test Image + Attribute Explanation prompt. This improvement stems primarily from its superior ability to parse geometry layers and read dimension labels. Crucially, its performance on reasoning-intensive parameters remains low. These findings underscore that orthographic projection CAD cannot be solved by prompting alone and requires dedicated fine-tuning strategy.
2) Prompt Format–Dependent Performance Gains. Appending an Attribute Explanation to the Test Image consistently boosts accuracy compared to the image-only baseline, with an average increase of 3 to 4 percentage points across models. This demonstrates that strong text encoders can leverage detailed, engineer-authored descriptions to guide complex geometric and numerical inferences. Similarly, providing a Reference Image or an Answered Pair results in comparable improvements, with a 4 to 5 percentage point increase in accuracy. These exemplars offer explicit visual-textual templates, simplifying the task of matching primitives to their corresponding semantic labels and dimensions. In contrast, using raw images forces models to address both perception and reasoning simultaneously, leading to the lowest performance. Collectively, these findings indicate that multimodal exemplars can enhance existing VLMs’ ability to reason over orthographic projection. However, this further underscores that additional fine-tuning strategies are crucial to fully bridging the remaining performance gap.
3)Introducing reasoning guidance yields consistent gains across all seven VLMs. Introducing reasoning guidance demonstrates that step-wise reasoning guidance helps models better decompose the multi-step orthographic projection tasks. In the absence of reasoning guidance, visual exemplar prompts (+Reference Image and +Answered Pair) deliver the largest relative improvements. However, when reasoning guidance is added, the Attribute Explanation prompt shows the greatest uplift. It stems from reasoning guidance is textual form, which synergizes most effectively with textual inputs. Hence, models with stronger text encoders (e.g., DeepSeek-V1) exhibit disproportionately larger boosts. These results underscore the necessity of integrating structured reasoning guidance to advance orthographic projection CAD reasoning.
Appendix C More performance on in-domain test set
Table 4: Performance comparison of training-free model, SFT model,and our GRPO-based model on orthographic projection reasoning tasks.
Prompts Without Reasoning Guidance Reasoning Guidance
Test Img +Reference +Answered +Attribute Test Img +Reference +Answered +Attribute
Img Pair Explanation Img Pair Explanation
Qwen2.5-VL 24.54 30.76 30.47 25.86 24.54 32.78 33.64 38.88
Qwen2.5-VL(SFT) 76.33 79.50 77.12 78.64 76.42 76.78 77.20 80.30
Ours 80.86 82.99 83.24 82.67 81.35 83.11 82.87 84.03
Tab. 4 presents the performance in-domain test set. The first row (Qwen2.5-VL) represents a training-free baseline, which exhibits consistently poor performance across all settings, indicating the inherent difficulty of this reasoning task without any fine-tuning or adaptation. The second row (Qwen2.5-VL with SFT) demonstrates a significant performance improvement, particularly when provided with reference images, answered pairs, and attribute explanations. However, despite these gains, the model still suffers from notable drops in more complex reasoning tasks, especially those involving composite parameter computation. In contrast, our method consistently outperforms both baselines across all settings, achieving the highest accuracy in all prompt configurations, with particularly strong results under reasoning guidance. These results validate the effectiveness of our approach in handling complex geometric reasoning and highlight its robustness across both simple and compositional inference scenarios.
Appendix D Failure cases and analysis
Figure 5: Four sets of failure cases. Each set consists of three parts: the left part shows the orthographic projection, the middle part presents the failed 3D model construction, and the right part illustrates the correctly constructed 3D model.
As illustrated in Fig. 5, we present four distinct failure cases, each highlighting different types of errors that can occur during the 3D model reconstruction process due to incorrect 2D parameterization. Fig. 5(a) demonstrates the error induced by incorrect dimensional parameters. Specifically, due to errors in the parameterization of certain dimensions, the constructed 3D model exhibits significant scaling issues, leading to a distorted and non-accurate representation. This failure emphasizes the critical importance of precise dimensional input when translating from 2D projections to 3D models, as even minor errors in size parameters can drastically affect the final output. Fig. 5(b) illustrates a failure caused by incorrect counting, which results in an incorrect number of primitives being identified and subsequently integrated into the 3D model. The discrepancy in the number of primitives reflects the direct impact that improper counting and recognition of geometric elements can have on the success of the 3D reconstruction. In Fig. 5(c), we observe a stacking issue caused by errors in the calculation of composite parameters. The failure manifests as improper alignment of components in the final model, where parts of the 3D structure fail to align correctly with each other. Lastly, Fig. 5(d) shows a failure resulting from exceeding physical boundaries due to incorrect composite parameter calculations. In this case, the incorrect parameters lead to elements of the 3D model extending beyond the physical constraints or limits of the original design. Together, these four failure cases provide strong evidence that errors in the parameterization of 2D models directly lead to inaccuracies in 3D model reconstruction. Each case illustrates the cascading effect of parameter errors from the 2D representation to the final 3D model, further underscoring the necessity for robust and precise parameterization methods in 2D-to-3D model conversion processes.

### v2

@@
 Abstract
-Computer-Aided Design (CAD) plays a pivotal role in industrial manufacturing.
-Orthographic projection reasoning underpins the entire CAD workflow, encompassing design, manufacturing, and simulation.
+Computer-Aided Design (CAD) is pivotal in industrial manufacturing, with orthographic projection reasoning foundational to its entire workflow—encompassing design, manufacturing, and simulation.
 However, prevailing deep-learning approaches employ standard 3D reconstruction pipelines as an alternative, which often introduce imprecise dimensions and limit the parametric editability required for CAD workflows.
 Recently, some researchers adopt vision–language models (VLMs), particularly supervised fine-tuning (SFT), to tackle CAD-related challenges.
-SFT shows promise but often devolves into pattern memorization, yielding poor out-of-distribution performance on complex reasoning tasks.
-To address these gaps, we introduce CReFT-CAD, a two-stage fine-tuning paradigm that first employs a curriculum-driven reinforcement learning stage with difficulty-aware rewards to build reasoning ability steadily, and then applies supervised post-tuning to hone instruction following and semantic extraction.
-Complementing this, we release TriView2CAD, the first large-scale, open-source benchmark for orthographic projection reasoning, comprising 200,000 synthetic and 3,000 real-world orthographic projections with precise dimension annotations and six interoperable data modalities.
-We benchmark leading VLMs on orthographic projection reasoning and demonstrate that CReFT-CAD substantially improves reasoning accuracy and out-of-distribution generalizability in real-world scenarios, offering valuable insights for advancing CAD reasoning research.
+SFT shows promise but often devolves into pattern memorization, resulting in poor out-of-distribution (OOD) performance on complex reasoning tasks.
+To tackle these limitations, we introduce CReFT-CAD, a two-stage fine-tuning paradigm: first, a curriculum-driven reinforcement learning stage with difficulty-aware rewards to steadily build reasoning abilities; second, supervised post-tuning to refine instruction following and semantic extraction.
+Complementing this, we release TriView2CAD, the first large-scale, open-source benchmark for orthographic projection reasoning, comprising 200,000 synthetic and 3,000 real-world orthographic projections with precise dimensional annotations and six interoperable data modalities.
+Benchmarking leading VLMs on orthographic projection reasoning, we show that CReFT-CAD significantly improves reasoning accuracy and OOD generalizability in real-world scenarios, providing valuable insights to advance CAD reasoning research.
 The code and adopted datasets are available at https://github.com/KeNiu042/CReFT-CAD.
@@
 1 Introduction
-Computer-Aided Design (CAD) is now integral to industrial product development, driving design, manufacturing, and simulation workflows.
-In the design phase, engineers use CAD drawings or rasterized orthographic projections for their precision and easy editability.
-During manufacturing, these drawings are converted into constraint-based parameter tables, and for simulation, they yield boundary-representation (B-Rep) data or textual geometry descriptions.
-A truly user-centric pipeline thus requires accurate semantic parsing of orthographic projection: automated extraction of parameter tables directly produces 3D models that meet stringent manufacturing and simulation standards.
-By contrast, existing reverse-engineering methods rely on expensive 2D/3D scanning hardware and labor-intensive post-processing, severely limiting scalability and broad industrial adoption.
-With the advent of deep learning, orthographic projection reasoning has typically been cast as a standard 3D reconstruction pipeline: rasterized drawings are processed by models that output B-Rep structures [1, 2], point clouds [3, 4], or meshes [5].
-However, two critical shortcomings remain.
-First, industrial drawings allow virtually zero tolerance for missing or erroneous components: pixel-level discretization errors in reconstruction can propagate into the 3D model, causing failures in downstream manufacturing and simulation.
-Second, practical CAD workflows demand parametric editability and precise semantic alignment—capabilities that common 3D representations lack.
-Although some methods generate CAD command sequences to restore editability, they still overlook the nuanced semantics of orthographic views and thus fail to map 2D drawing features to their intended 3D counterparts reliably.
-Recent works leverage the modality alignment and semantic reasoning strengths of vision–language models (VLMs) [6, 7, 8, 9, 10, 11], applying supervised fine-tuning (SFT) to address CAD challenges[12, 13, 14].
-However, orthographic projection reasoning demands genuine inferential ability, and task-specific SFT often leads to memorization and out-of-distribution (OOD) degradation rather than deeper reasoning [15].
-Inspired by DeepSeek R1-Zero’s Group Relative Policy Optimization (GRPO) [16], we propose Curriculum-driven Reinforcement Fine-tuning for CAD (CReFT-CAD),
-CReFT-CAD is a versatile framework that assists engineers in achieving orthographic projection reasoning throughout the entire lifecycle of industrial design, from design and manufacturing to simulation, in an interactive manner, as shown in Fig.
-1.
-CReFT-CAD is a two-stage paradigm that combines reinforcement learning and supervised tuning.
-Curriculum-driven Reinforcement Fine-tuning: We introduce a difficulty-aware reward mechanism that incrementally escalates task complexity, gradually exposing the model to more complex reasoning challenges and promoting stable policy optimization.
-Supervised Post-tuning: Building on the reinforced model, we further refine its instruction-following and reasoning capabilities via multi-task SFT.
-Our qualitative and quantitative evaluations demonstrate that CReFT-CAD delivers robust orthographic projection reasoning and strong generalization in OOD scenarios.
-Another major obstacle in orthographic projection reasoning has been the lack of open-source datasets with high-fidelity annotations.
-To fill this gap, we introduce TriView2CAD, the first large-scale benchmark tailored for industrial CAD pipelines.
-Leveraging real-world design archives, we employ a constraint-driven synthesis pipeline to generate 200,000 synthetic and 3,000 real-world three-view sets, each annotated with precise, one-to-one dimension labels tied to their corresponding geometric primitives.
-This design enables models to extract quantitative measurements from rasterized drawings and enforce exact geometric constraints for simulation-ready 3D reconstruction.
+Computer‐Aided Design (CAD) is now integral to industrial product development, underpinning design, manufacturing, and simulation workflows.
+In the design phase, engineers use CAD drawings or rasterized orthographic projections for their precision and facile editability.
+During manufacturing, these drawings are converted into constraint‐based parameter tables; for simulation, they yield boundary‐representation (B‐Rep) data or textual geometry descriptions.
+A truly user‐centric pipeline thus requires accurate semantic parsing of orthographic projections: automated extraction of parameter tables enables direct generation of 3D models meeting stringent manufacturing and simulation standards.
+By contrast, existing reverse‐engineering methods rely on expensive 2D/3D scanning hardware and labor‐intensive post-processing, greatly restricting scalability and widespread industrial adoption.
+With the advent of deep learning, orthographic projection reasoning has typically been framed as a standard 3D reconstruction pipeline: models process rasterized drawings to output B-Rep structures [1, 2], point clouds [3, 4], or meshes [5].
+However, this paradigm suffers from two critical limitations.
+First, industrial drawings demand near-zero tolerance for missing or erroneous components: pixel-level discretization errors in reconstruction can propagate to 3D models, leading to failures in downstream manufacturing and simulation.
+Second, practical CAD workflows require parametric editability and precise semantic alignment—key capabilities that common 3D representations lack.
+While some methods generate CAD command sequences to recover editability, they still overlook the nuanced semantics of orthographic views, thus failing to reliably map 2D drawing features to their intended 3D counterparts.
+Recent works leverage the modality alignment and semantic reasoning capabilities of vision–language models (VLMs) [6, 7, 8, 9, 10, 11, 12], addressing CAD challenges via supervised fine-tuning (SFT) [13, 14, 15].
+However, orthographic projection reasoning requires genuine inferential ability, and task-specific SFT often induces pattern memorization and out-of-distribution (OOD) performance degradation rather than fostering deeper reasoning [16].
+Inspired by DeepSeek R1-Zero’s Group Relative Policy Optimization (GRPO) [17], we propose Curriculum-driven Reinforcement Fine-tuning for CAD (CReFT-CAD).
+CReFT-CAD is a versatile framework enabling interactive orthographic projection reasoning across the full lifecycle of industrial design—from design and manufacturing to simulation (see Fig.
+1).
+It consists of a two-stage paradigm integrating reinforcement learning and supervised tuning: the first stage is Curriculum-driven Reinforcement Fine-tuning, where we introduce a difficulty-aware reward mechanism that incrementally increases task complexity, gradually exposing the model to increasingly complex reasoning challenges to promote stable policy optimization; the second stage is Supervised Post-tuning, where we build on the reinforced model and further refine its instruction-following and reasoning capabilities via multi-task SFT.
+Our qualitative and quantitative evaluations show that CReFT-CAD achieves robust orthographic projection reasoning and strong generalization in OOD scenarios.
+Another key barrier in orthographic projection reasoning is the scarcity of open-source datasets with high-fidelity annotations.
+To address this, we present TriView2CAD, the first large-scale benchmark tailored to industrial CAD pipelines.
+Leveraging real-world design archives, we use a constraint-driven synthesis pipeline to generate 200,000 synthetic and 3,000 real-world three-view sets, each annotated with precise, one-to-one dimensional labels linked to their corresponding geometric primitives.
+This design allows models to extract quantitative measurements from rasterized drawings and enforce exact geometric constraints for 3D reconstructions ready for simulation.
 Figure 1: (a) (b) show the results of parameterization tasks of orthographic projection reasoning using Qwen2.5-VL and Deepseek without tuning.
@@
 Understanding and reasoning with orthographic projection is a fundamental and longstanding challenge in CAD-related tasks [18, 19, 20].
-With the development of 3D reconstruction methods, existing orthographic projection reasoning methods have largely focused on standard 3D reconstruction pipelines.
-SPARE3D [20] introduces a dataset designed to evaluate the spatial reasoning capabilities of AI systems through 2D line drawings of 3D objects.
-Contrastive-SPARE3D [21] propose a self-supervised binary classification network that helps to learn 3D object line drawings representations that are detail-sensitive and view-invariant.
-PlankAssembly [22] introduces a transformer-based sequence generation model that learns flexible mappings between inputs and outputs.
-IsoTGAN [23] introduces a novel Gaussian-enhanced Euclidean attention mechanism and a geometric constraint loss function to enhance local image features further.
-GaussianCAD [24] employs a custom sparse-view 3D reconstruction method, eliminating the reliance on vector CAD sketches and 3D real-world data.
+With the development of 3D reconstruction methods, existing methods for orthographic projection reasoning have largely centered on standard 3D reconstruction pipelines.
+SPARE3D [21] introduces a dataset for evaluating the spatial reasoning capabilities of AI systems via 2D line drawings of 3D objects.
+Contrastive-SPARE3D [22] proposes a self-supervised binary classification network that helps learn 3D object line drawing representations that are detail-sensitive and view-invariant.
+PlankAssembly [23] presents a transformer-based sequence generation model that learns flexible input-output mappings.
+IsoTGAN [24] introduces a novel Gaussian-enhanced Euclidean attention mechanism and a geometric constraint loss function to further enhance local image features.
+GaussianCAD [25] employs a custom sparse-view 3D reconstruction method, removing reliance on vector CAD sketches and real-world 3D data.
 2.2 Applications of LVLMs in CAD-Related tasks
-Recent work on applying vision–language models (VLMs) to CAD tasks can be grouped into three main categories: 1) Direct Adaptation of Pretrained VLMs.
-CAD-Recode [25] leverages a VLM to translate raw point-cloud inputs into executable Python code.
-Both LEAM [26] and LLM4CAD [27] generate CAD models by jointly leveraging textual descriptions and image inputs.
-CAD-Assistant [28] introduces a novel approach that leverages VLMs as planners, where the model interacts with Python APIs to accomplish various CAD tasks.
-2) Supervised Fine-Tuning (SFT).
-CAD2Program [12] generates 3D parametric models from 2D CAD drawings.
-CAD-MLLM [29] enables parametric CAD modeling from text, images, and point clouds, and proposes the Omni-CAD dataset.
-3) Reinforcement Learning–Based Tunings.
-RLCAD [30] proposes a reinforcement learning environment to generate complex CAD models.
-CADCrafter [31] presents an image-to-CAD model generation framework, fine-tuned via Direct Preference Optimization (DPO) to translate input images directly into executable CAD representations.
+Recent efforts applying vision–language models (VLMs) to CAD tasks fall into three main categories:1) Direct Adaptation of Pretrained VLMs.
+CAD-Recode [26] leverages a VLM to translate raw point cloud inputs into executable Python code.Both LEAM [27] and LLM4CAD [28] generate CAD models by jointly leveraging text descriptions and image inputs.
+CAD-Assistant [29] introduces a novel approach using VLMs as planners, where the model interacts with Python APIs to accomplish diverse CAD tasks.
+2) Supervised Fine-Tuning (SFT).CAD2Program [13] generates 3D parametric models from 2D CAD drawings.
+CAD-MLLM [30] enables parametric CAD modeling from text, images, and point clouds, and introduces the Omni-CAD dataset.
+3) Reinforcement Learning–based Tunings.RLCAD [31] introduces a reinforcement learning environment for generating complex CAD models.
+CADCrafter [32] presents an image-to-CAD generation framework, fine-tuned via Direct Preference Optimization (DPO) to directly translate input images into executable CAD representations.
 3 TriView2CAD: Benchmarking orthographic projection reasoning
@@
 Existing orthographic projection benchmarks provide only rasterized drawings without precise dimension annotations.
-It causes models to prioritize relative scale over exact measurements, and pixel-level errors in raster images can easily propagate into the reconstructed 3D model, resulting in inaccuracies.
-Additionally, the limited editability of reconstructed 3D models precludes corrective adjustments during design and simulation.
+This causes models to prioritize relative scale over exact measurements, and pixel-level errors in rasterized images can easily propagate into the reconstructed 3D model, resulting in inaccuracies.
+Additionally, the limited editability of reconstructed 3D models prevents corrective adjustments during design and simulation.
 Lack of explicit logical reasoning tasks.
-No current dataset evaluates a model’s inferential capabilities within or across orthographic views.
-Consequently, models trained on these benchmarks learn pixel-level reconstruction but fail to handle real-world challenges such as inferring omitted annotations or exploiting structural symmetries.
+Existing datasets do not evaluate a model’s inferential capabilities within or across orthographic views.
+Consequently, models trained on these benchmarks learn pixel-level reconstruction skills but fail to handle real-world challenges such as inferring omitted annotations or leveraging structural symmetries.
 Lack of essential CAD modalities.
-Most datasets include only one or two modalities—typically raster images and, at best, a single 3D format (meshes or point clouds).
-This limits coverage omits critical data representations (e.g., parameter tables, vector CAD files, executable commands, STEP/B-Rep) needed for end-to-end CAD design, manufacturing, and simulation pipelines.
+Most datasets contain only one or two modalities—typically raster images and, at best, a single 3D format (meshes or point clouds).
+This limited coverage omits critical data representations (e.g., parameter tables, vector CAD files, executable commands, STEP/B-Rep) required for end-to-end CAD design, manufacturing, and simulation pipelines.
 3.1 Dataset construction
-In this work, we focus on prefabricated bridge piers due to their inherently modular composition of repeated, standardized components, which yields a compact yet expressive parameter space for dataset synthesis.
-Moreover, prefabricated piers are ubiquitous in modern infrastructure and supported by extensive archives of CAD drawings and manufacturing documentation.
+In this work, we focus on prefabricated bridge piers owing to their inherently modular structure—composed of repeated, standardized components—that enables a compact yet expressive parameter space for dataset synthesis.
+Moreover, prefabricated piers are ubiquitous in modern infrastructure and backed by extensive CAD drawing and manufacturing documentation archives.
 TriView2CAD comprises 200,000 synthetic samples and 3,000 real-world samples.
-The synthetic data are split 80/20 into training and test sets, while, due to the scarcity of real-world engineering drawings, all 3,000 real-world samples are reserved exclusively for testing.
-The synthetic test set is used to evaluate in-domain performance.
+The synthetic data are split 80/20 into training and test sets; due to the scarcity of real-world engineering drawings, all 3,000 real-world samples are reserved exclusively for testing.
+The synthetic test set evaluates in-domain performance.
 As shown in Fig.
-3, real-world images typically exhibit more complex structures, primarily due to redundant annotation lines causing occlusions, overlap with other components, and the inclusion of non-numeric annotations, serving to assess out-of-distribution generalization in authentic CAD scenarios.
-To mirror end-to-end industrial workflows, TriView2CAD provides six interconnected data modalities per sample—structured parameter tables (JSON), vector CAD drawings (DXF), raster images (PNG), executable modeling scripts, and two standard 3D formats (STEP and B-Rep).
-This comprehensive suite supports tasks from early design to downstream manufacturing and simulation.
+3, real-world images typically feature more complex structures—primarily due to redundant annotation lines (causing occlusions and overlaps with other components) and the inclusion of non-numeric annotations.
+These samples thus assess out-of-distribution (OOD) generalization in authentic CAD scenarios.
+To align with end-to-end industrial workflows, TriView2CAD provides six interconnected data modalities per sample: structured parameter tables (JSON), vector CAD drawings (DXF), raster images (PNG), executable modeling scripts, and two standard 3D formats (STEP and B-Rep).
+This comprehensive modality suite supports tasks spanning early design to downstream manufacturing and simulation.
 As shown in Fig.
-5, we employ a constraint-driven synthesis pipeline.
-The detailed steps are as follows:
+5, we adopt a constraint-driven synthesis pipeline, with detailed steps outlined below:
 Figure 2: Constraint-driven synthesis pipeline for TriView2CAD.
 Figure 3: Examples of Real-World Orthographic Projections and its 3D model.
-Step 1: Constraint-Guided Parameter Sampling We first analyze real-world CAD drawings to define a 15-dimensional parameter space subject to two constraint classes.
-Intra-View Constraints enforce both topological closure and physical validity within each projection: every component must form a gap- and overlap-free contour, and paired dimensions must satisfy domain-specific engineering constraints (e.g., “Cross-Bridge Pier Spacing” < “Cap Beam Cross-Bridge Dimension”).
-Inter-View Constraints ensure cross-projection consistency by requiring every measurement annotated in one view to exactly match its counterpart (height, width, and depth) in the other views.
-By embedding these constraints into our sampling pipeline, we guarantee that every synthetic design is geometrically coherent and adheres to real-world engineering practices.
-Step 2: 2D Drawing and Raster Image Generation We employ the ezdxf library to translate each sampled parameter vector into a vectorized 2D CAD drawing (DXF), mapping numeric dimensions to primitives such as lines, circles, and arcs to form coherent orthographic projections.
-We leverage ezdxf’s layers-separating function to partition primitives into semantic layers, thereby enhancing the editability of the resulting CAD drawings.
-Each DXF file is then loaded into FreeCAD [32], where we capture high-resolution screenshots to produce three orthographic views (front, top, and side).
-Step 3: 3D Model Generation We employ the FreeCAD Python API to reconstruct 3D model from its parameter vector.
-Mirroring our 2D pipeline, we decompose the model into individual entities and generate scripted commands that accurately position each primitive according to the intra- and inter-view constraints established in Step 1.
-By executing these scripts, we automatically produce industry-standard STEP and B-Rep files, closing the loop from parameter sampling to editable, simulation-ready CAD geometry.
+Step 1: Constraint-Guided Parameter Sampling We first analyze real-world CAD drawings to define a 15-dimensional parameter space, governed by two constraint classes.
+Intra-View Constraints enforce topological closure and physical validity within each projection: every component must form gap- and overlap-free contours, and paired dimensions adhere to domain-specific engineering constraints (e.g., “Cross-Bridge Pier Spacing” < “Cap Beam Cross-Bridge Dimension”).
+Inter-View Constraints ensure cross-projection consistency by requiring measurements annotated in one view to exactly match their counterparts (height, width, depth) in other views.
+By embedding these constraints into the sampling pipeline, we ensure all synthetic designs are geometrically coherent and compliant with real-world engineering standards.
+Step 2: 2D Drawing and Raster Image Generation We use the ezdxf library to convert each sampled parameter vector into a vectorized 2D CAD drawing (DXF format), mapping numeric dimensions to geometric primitives (e.g., lines, circles, arcs) to form coherent orthographic projections.
+We utilize ezdxf’s layer separation functionality to partition primitives into semantic layers, enhancing the editability of the resulting CAD drawings.
+Each DXF file is then imported into FreeCAD [33], where high-resolution screenshots are captured to generate three orthographic views (front, top, side).
+Step 3: 3D Model Generation We utilize the FreeCAD Python API to reconstruct a 3D model from its corresponding parameter vector.
+Consistent with our 2D generation pipeline, we decompose the model into discrete entities and generate scripted commands to precisely position each primitive in compliance with the intra- and inter-view constraints defined in Step 1.
+Executing these scripts automatically produces industry-standard STEP and B-Rep files, closing the loop from parameter sampling to editable, simulation-ready CAD geometries.
 3.2 Evaluation of TriView2CAD
-We benchmark seven leading vision–language models on TriView2CAD to perform high-precision orthographic projection reasoning.
+We benchmark seven leading vision–language models on TriView2CAD to assess their high-precision orthographic projection reasoning capabilities.
 The evaluation results are presented in Sec.
 3.2.
-Reflecting real-world CAD workflows, we define three complementary evaluation tasks to rigorously test a model’s ability to extract precise dimension annotations from rasterized orthographic projection and infer the underlying geometric relationships.
-(i) Dimension recognition and pairing Models identify each annotated dimension in a single orthographic view and map it to its corresponding geometric feature.
-(ii) Primitive counting Models count instances of specified CAD elements (e.g., number of pier columns), testing their ability to parse structural composition.
-(iii) Composite Parameter Computation Models compute engineering-critical derived quantities—for example, the Cross-Bridge Pier Spacing is defined as the sum of the Pier Column Cross-Bridge Dimension and the Pile Spacing.
-Overall, we define a 15-dimensional parameter space comprising 6 recognition parameters, 3 counting parameters, and 6 composite calculation parameters.
-During evaluation, we treat each individual parameter as a separate prediction.
-The overall accuracy is then computed as the total number of correctly predicted parameters divided by the total number of parameters across all test samples.
-This evaluation suite exposes each model’s strengths and weaknesses in handling the tightly coupled semantic and quantitative demands of orthographic projection CAD reasoning.
+Aligning with real-world CAD workflows, we design three complementary evaluation tasks to rigorously test a model’s ability to extract precise dimension annotations from rasterized orthographic projections and infer the underlying geometric relationships: (i) Dimension recognition and pairing Models identify each annotated dimension in a single orthographic view and map it to its associated geometric feature.
+(ii) Primitive counting Models count instances of specified CAD elements (e.g., number of pier columns), assessing their ability to parse structural composition.
+(iii) Composite Parameter Computation Models compute engineering-critical derived quantities—for example, Cross-Bridge Pier Spacing is calculated as the sum of Pier Column Cross-Bridge Dimension and Pile Spacing.
+Collectively, we define a 15-dimensional parameter space consisting of 6 recognition parameters, 3 counting parameters, and 6 composite calculation parameters.
+In evaluation, each parameter is treated as an independent prediction target.
+Overall accuracy is computed as the total number of correctly predicted parameters divided by the total number of parameters across all test samples.
+This evaluation suite reveals the strengths and weaknesses of each model in addressing the tightly coupled semantic and quantitative requirements of orthographic projection CAD reasoning.
 4 Methodology
-Inspired by the success of reinforcement learning–based tuning methods [33, 34], we present Progressive Reinforcement Fine-tuning for CAD (CReFT-CAD), a orthographic-projection reasoning framework.
-CReFT-CAD combines a ViT-based visual encoder with Qwen2.5 [35].
-First, Progressive Reinforcement Fine-tuning employs a curriculum of three progressively harder reasoning tasks under a difficulty-aware reward scheme, enabling the model to move beyond rote pattern matching toward robust reasoning while improving OOD resilience.
-Second, Supervised Post-tuning refines the model’s instruction-following to support interactive, real-world CAD queries.
-4.1 Curriculum-driven reinforcement fine-tuning
-Orthographic projection reasoning is a complex reasoning task that encompasses not only dimension recognition and pairing but also primitive counting and more intricate composite parameter computation tasks.
-Building on GRPO, we propose a Curriculum-driven Reinforcement Fine-tuning (CReFT) strategy, which incrementally exposes the model to progressively more complex training tasks to enhance its reasoning and generalization capabilities.
-Our tasks design mirrors the core steps of expert engineers in verifying CAD drawings, incorporating dichotomous choice task, multiple choice task, and parameterization tasks based on chain-of-thought (CoT).
-For each task, we design a tailored, efficient reward function to ensure the model can steadily and stably optimize its reasoning performance.
-Figure 4: Diagram design of three tasks in curriculum-driven reinforcement fine-tuning.
+Inspired by the success of reinforcement learning–based tuning methods [34, 35], we present Curriculum-driven Reinforcement Fine-tuning for CAD (CReFT-CAD), an orthographic-projection reasoning framework.
+CReFT-CAD integrates a ViT-based visual encoder with Qwen2.5 [36].
+First, the framework adopts a curriculum of three progressively challenging reasoning tasks paired with a difficulty-aware reward mechanism, empowering the model to transcend rote pattern matching and develop robust reasoning capabilities while enhancing out-of-distribution (OOD) generalization.
+Second, Supervised Post-tuning refines the model’s instruction-following abilities to accommodate interactive, real-world CAD queries.
+4.1 Curriculum‐driven reinforcement fine-tuning
+Orthographic projection reasoning is a sophisticated reasoning task that involves not only dimension recognition and pairing but also primitive counting and more complex composite parameter computation.
+Building upon GRPO, we propose a Curriculum-driven Reinforcement Fine-tuning (CReFT) strategy, which incrementally exposes the model to increasingly complex training tasks to enhance its reasoning and generalization capabilities.
+Our task design aligns with the core steps expert engineers follow when verifying CAD drawings, incorporating dichotomous choice tasks, multiple-choice tasks, and chain-of-thought (CoT)-based parameterization tasks.
+For each task, we design a tailored, efficient reward function to ensure the model can steadily optimize its reasoning capabilities.
+Figure 4: Diagram design of three tasks in curriculum‐driven reinforcement fine-tuning.
 Data engine.
-We leverage 160,000 image–text instruction pairs using the synthetic training set of TriView2CAD, where each prompt includes a given orthographic projection and all 15 parameter key-value pairs, as shown in Fig.
-4.
-To ensure a balanced training signal across all three tasks, 50% of the training responses are entirely correct, while the remaining 50% contain errors of varying magnitude.
-Specifically, for the dichotomous choice task, we ensure that each negative response contains $n$ erroneous parameter values, where $n$ follows a uniform distribution $N\sim\text{Uniform}(a,b)$, where $a=1$ and $b=15$.
-For the multiple choice task, $p$ parameter values are masked in each image-text instruction pair.
-The masking mechanism ensures the diversity of options.
-This approach prevents the model from simply deriving the answer through parameterization, bypassing the complex multiple choice reasoning required in the task.
-Instead, it forces the model to engage in decision-making actively, considering multiple potential solutions and selecting the correct responses based on a deeper understanding.
-For each instruction pair, unmasked values are correct, while for the incorrect parameter value lists, all unmasked values, except for randomly selected $q$ erroneous lists.
-Both the masked values $p$ and the number of incorrect values $q$ follow a normal distribution, introducing a smooth and gradual variation in the difficulty of the learning task.
-This variation helps the model to learn in a more structured and progressive manner.
-For the parameterization task, we construct image-text instruction pairs based on Chain-of-Thought (CoT) reasoning.
-Specifically, as for Composite Parameter Computation, expert knowledge guides the model through a step-by-step reasoning guidance to arrive at the correct output.
-The step-by-step workflow of reasoning guidance is as follow: 1) recognize parameters required for the computation of the composite parameter; 2) provide the calculation formula for the composite parameter; 3) generate the predicted result based on the formula and the factor parameters
+We utilize 160,000 image–text instruction pairs from the synthetic training set of TriView2CAD, where each prompt includes an orthographic projection and all 15 parameter key-value pairs (see Fig.
+4).To ensure a balanced training signal across the three tasks, 50% of the training responses are fully correct, while the remaining 50% contain errors of varying degrees.
+Specifically, for dichotomous choice tasks, each negative response includes $n$ erroneous parameter values, where $n$ follows a uniform distribution $N\sim\text{Uniform}(a,b)$, where $a=1$ and $b=15$.
+For multiple-choice tasks, $p$ parameter values are masked out in each image-text instruction pair.
+The masking mechanism is designed to ensure the diversity of distractor options, preventing the model from merely deriving answers through parameter matching—thereby avoiding circumvention of the complex reasoning required for the task.
+Instead, it compels the model to actively engage in decision-making, evaluating multiple potential solutions and selecting the correct one based on a deeper geometric understanding.
+For each instruction pair, unmasked values are correct; for incorrect parameter value lists, all unmasked values are correct except for those in $q$ randomly selected erroneous entries.
+Both the number of masked values $p$ and the number of erroneous entries $q$ follow normal distributions, introducing smooth, gradual variations in task difficulty that enable the model to learn in a more structured, progressive fashion.
+For parameterization tasks, we construct image-text instruction pairs grounded in Chain-of-Thought (CoT) reasoning.
+Specifically, for Composite Parameter Computation tasks, expert knowledge is embedded into step-by-step reasoning prompts to guide the model toward correct outputs.
+The step-by-step reasoning workflow is as follows: 1) identify the parameters required to compute the composite parameter; 2) specify the calculation formula for the composite parameter; 3) compute the predicted result using the formula and relevant factor parameters.
 Training strategy.
-Given that our task involves highly customized inputs and outputs rather than a general-purpose problem, directly applying a pretrained VLM for reasoning tasks and generating required outputs proves ineffective.
-As a result, the model cannot effectively execute the reinforcement learning reward mechanism, as the reward will always be 0.
-To address this, prior to executing each training task, we pre-sample a subset of image-text instruction pairs to "warm up" the model, facilitating the initialization of the reinforcement learning (RL) reward mechanism.
+Given that our task involves highly specialized inputs and outputs (as opposed to general-purpose problems), directly applying a pretrained VLM to perform reasoning and generate the required outputs is ineffective.
+Consequently, the reinforcement learning (RL) reward mechanism cannot function properly—with rewards consistently remaining zero.
+To mitigate this, we pre-sample a subset of image-text instruction pairs to "warm up" the model before initiating each training task, enabling proper initialization of the RL reward mechanism.
 4.1.1 Task 1: Dichotomous choice task
-As for this task, the model’s output is constrained to dichotomous responses, either "yes" or "no." A "yes" answer is given only when the entire 15-dimensional parameter space is completely correct, while any error, even in a single parameter, results in a "no" response.
-This strict criterion ensures that the model is incentivized to produce fully accurate parameterizations.
-To align with this setup, the reward function is defined as follows:
+For this task, the model’s output is restricted to dichotomous responses: "yes" or "no." A "yes" response is returned if and only if all parameters in the 15-dimensional parameter space are fully accurate, whereas any error—even in a single parameter—triggers a "no" response.
+This stringent criterion incentivizes the model to generate fully accurate parameterizations.
+Consistent with this design, the reward function is defined as follows:
 $R_{\text{P1}}=\begin{cases}1,&\text{ if all parameters are correct}\\ 0,&\text{otherwise}\end{cases}$ (1)
@@
 Training strategy.
-In the task 2, we introduce multiple choice tasks to further increase the complexity of decision-making.
-These tasks require the model to select from multiple parameter value lists, which significantly enhancing the cognitive load compared to binary decisions.
-Building on the first stage, we pre-sample a subset of image-text instruction pairs to perform instruction-following training on the LLM.
-This training allows the model to generate structured outputs in the form of multiple choice questions, enabling the reinforcement learning reward mechanism.
-Additionally, we designed a stringent reward function that encourages the model to maximize the utilization of image-text instruction pairs, ensuring that it extracts and applies all relevant information from both the visual and textual components.
-The reward function incentivizes the model to select the most accurate option from the provided set of choices.
-Any incorrect selection results in a reward of 0, thereby pushing the model to focus on precision and reliability in its decision-making process.
+Task 2 introduces multiple-choice tasks to further increase decision-making complexity.
+These tasks require the model to select from multiple parameter value lists, which significantly enhances the cognitive load compared to binary decisions.
+Building on the first task, we pre-sample a subset of image-text instruction pairs to conduct instruction-following training on the LLM.
+This training enables the model to generate structured outputs in the form of multiple-choice questions, facilitating the activation of the reinforcement learning (RL) reward mechanism.
+Additionally, we design a stringent reward function that encourages the model to fully utilize image-text instruction pairs, ensuring it extracts and leverages all relevant information from both visual and textual modalities.
+The reward function incentivizes the model to select the most accurate option from the provided choices.
+Any incorrect selection yields a reward of 0, compelling the model to prioritize precision and reliability in its decision-making.
 Let $S_{\text{correct}}$ be the set of correct parameter value lists, $S_{\text{incorrect}}$ be the set of incorrect parameter value lists, and $S_{\text{Selected}}$ be the set of selected parameter value lists by the model.
 The reward function $R_{p2}$ is defined as:
-$R_{\text{P2}}=\begin{cases}1&\text{if }S_{\text{selected}}=S_{\text{correct}}\\ 0.2&\text{if }S_{\text{selected}}\supseteq S_{\text{correct}}\text{ but }S_{\text{selected}}\neq S_{\text{correct}}\\ 0&\text{if }(S_{\text{selected}}\cap S_{\text{correct}}\neq\emptyset)\text{ and }(S_{\text{selected}}\cap S_{\text{incorrect}}\neq\emptyset)\\ 0&\text{if }S_{\text{selected}}=S_{\text{incorrect}}\end{cases}$ (2)
-Where $S_{\text{Selected}}=S_{\text{correct}}$ means the model selects only the correct parameter value lists; $S_{\text{Selected}}\supseteq S_{\text{correct}}\text{ but }S_{\text{Selected}}\neq S_{\text{correct}}$ means the model selects all correct lists but fails to select all of them; $S_{\text{Selected}}\cap S_{\text{correct}}\neq\emptyset$ and $S_{\text{Selected}}\cap S_{\text{incorrect}}\neq\emptyset$ means the model selects both correct and incorrect parameter value lists; $S_{\text{Selected}}=S_{\text{incorrect}}$ means the model selects incorrect parameter value lists.
+[…]
+[…]
 4.1.3 Task 3: Parameterization tasks based on chain-of-thought
 Training strategy.
-In the training process, we begin with the model trained in task 2, and sample synthetic data to construct standard orthographic projection reasoning question-answer pairs.
-These pairs are then used to perform small-batch fine-tuning.
-Once the model is fine-tuned, we assess the difficulty of predicting the 15-dimensional parameters by evaluating the model’s performance on a test set.
-The prediction accuracy for each parameter is categorized into three difficulty levels: Easy Attributes: Attributes with an accuracy greater than 0.8.
+For this task’s training, we start with the model fine-tuned in Task 2 and sample synthetic data to construct standard orthographic projection reasoning QA pairs.
+These pairs are used for small-batch fine-tuning.
+After fine-tuning, we assess the difficulty of predicting the 15-dimensional parameters by evaluating the model’s performance on a held-out test set.
+Each parameter’s prediction accuracy is categorized into three difficulty levels: Easy Attributes: Attributes with an accuracy greater than 0.8.
 Medium Attributes: Attributes with accuracy between 0.2 and 0.8.
 Difficult Attributes: Attributes with accuracy less than 0.2.
-This difficulty classification is then used to design the reward function The reward for each correctly answered attribute is determined by its difficulty level as follows:
-$R_{\text{P3}}=\begin{cases}1&\text{if the model correctly predicts an easy attribute }\\ 1.5&\text{if the model correctly predicts a medium attribute }\\ 2&\text{if the model correctly predicts a hard attribute }\\ 0&\text{if the model's prediction is incorrect }\end{cases}$ (3)
+This difficulty classification is subsequently used to design the reward function.
+The reward for each correctly predicted attribute is determined by its difficulty level as follows:
+[…]
 This reward function ensures that the model is motivated to gradually improve across all difficulty levels, with greater emphasis on mastering the more challenging tasks, thus enhancing the model’s overall reasoning capabilities in orthographic projection tasks.
 4.2 Supervised post-tuning
-After the CReFT stage, where the model developed robust reasoning and generalization capabilities, the focus of the Supervised Post-tuning task shifts to enhancing the model’s instruction-following ability.
-During this stage, we rethink the industrial design application context and, with a user-centric approach, reformulate tasks in the style of Visual Question Answering (VQA).
-By leveraging multimodal language models, we design specific tasks that guide the model to output a comprehensive list of parameters and make critical judgments based on orthographic projection.
-In this stage, the model is trained to process the following tasks: 1) Full Parameter List Output Task: The goal is to generate a fully accurate set of parameter key-value pairs, which is the most common format for outputs during the simulation phase.
-2) Parameter Validation Task: The task is to compare the extracted parameter key-value pairs with the accurate parameter key-value pairs in the orthographic projection to determine whether the extracted parameters are correct.
-3) View Matching Task: The model is given two individual orthographic projection views.
-The model must, based on parameter constraints and rasterized images, assess whether the views correspond to the same overall 3D object or design.
-4) Component Counting Task: The model is required to identify and count specific components in the drawing, such as bridge piers or columns, based on the given parameter constraints.
-The output of this task helps to quantify the number of elements in the design, aiding in material estimation and component verification.
-By employing this post-tuning phase, we enhance the model’s ability to interpret and respond to complex design queries, thereby improving its interactive reasoning abilities and making it more applicable to real-world industrial design workflows.
+Following the CReFT stage—during which the model has developed robust reasoning and generalization capabilities—the focus of Supervised Post-tuning shifts to refining the model’s instruction-following capabilities.
+We revisit the industrial design application context and, adopting a user-centric approach, reformulate tasks in the paradigm of Visual Question Answering (VQA).
+Leveraging multimodal language models, we design specific tasks that guide the model to output comprehensive parameter lists and make engineering-critical judgments based on orthographic projections.
+In this stage, the model is trained to handle the following tasks: 1) Full Parameter List Output Task: Generate a complete and accurate set of parameter key-value pairs— the standard output format for the simulation phase.
+2) Parameter Validation Task: Compare extracted parameter key-value pairs with the ground-truth parameter set in the orthographic projections to verify the accuracy of the extracted parameters.
+3) View Matching Task: Given two distinct orthographic views, determine if the views correspond to the same 3D object or design based on parameter constraints and rasterized images.
+4) Component Counting Task: Identify and count specified components (e.g., bridge piers, columns) in the drawings based on given parameter constraints.
+This task quantifies design elements to support material estimation and component verification.
+Through this post-tuning phase, we strengthen the model’s capacity to interpret and respond to complex design queries, thereby enhancing its interactive reasoning capabilities and increasing its applicability to real-world industrial design workflows.
 5 Experiments
 Implementation Details.
-All experiments are conducted on NVIDIA A100 GPUs.
-Most experiments use GOT-OCR2.0 as the base model, trained on a single server equipped with 8 A100 GPUs and a batch size of 64.
-We use the AdamW optimizer with a cosine annealing scheduler.
-The hyperparameters are configured as follows: (1) Learning rate: 1e-6 for RL (GRPO) training and 2e-5 for baseline SFT experiments.
-(2) Maximum input image size: 2,483,776 pixels.
-(3) Total GRPO training steps: 1500.
+All experiments were conducted on NVIDIA A100 GPUs.
+Most experiments used GOT-OCR2.0 as the base model, trained on a single server with 8 A100 GPUs and a batch size of 64.
+We adopted the AdamW optimizer with a cosine annealing scheduler.
+The hyperparameters were configured as follows:(1) Learning rate: 1e-6 for RL (GRPO) training and 2e-5 for baseline SFT experiments;(2) Maximum input image size: 2,483,776 pixels;(3) Total GRPO training steps: 1500.
 Datasets and Metrics.
-We evaluate both the leading vision–language models (VLMs) and the proposed method on TriView2CAD, following the evaluation strategy outlined in Sec.
+We evaluated both leading vision–language models (VLMs) and our proposed method on TriView2CAD, following the evaluation strategy outlined in Sec.
 3.2.
-The synthetic test set is used to assess in-domain accuracy.
-Due to the difficulty in obtaining real-world data, all 3,000 real-world samples are reserved exclusively for testing, allowing us to evaluate out-of-distribution (OOD) generalization in real-world scenarios.
-We design four distinct prompt configurations to probe their orthographic projection reasoning capabilities: Test image-Only: Models receive only the target rasterized drawings.
-Test image + Reference Image: In addition to the target drawings, models are given a reference image—rendered on the same geometry layer—with all dimension names to supply semantic alignment.
-Test image + Answered Pair: Inputs consist of the target drawings plus a correctly paired raster image and its parameter vector, providing an exemplar mapping between visual features and quantitative parameters.
-Test image + Atttribute Explanation: Models are supplied with the target drawings alongside a detailed textual interpretation authored by professional engineers.
-Given that orthographic projection reasoning involves complex multi-step inference and current pretrained VLMs limit exposure to CAD drawings, we incorporate explicit reasoning guidance into each prompt format to encourage clear reasoning trajectories.
-For each of the four different prompt input formats, we outline detailed step-by-step instructions for completing the task.
-These steps are designed to guide the model through the reasoning process in a structured manner.
+The synthetic test set was used to assess in-domain accuracy.
+Due to the scarcity of real-world data, all 3,000 real-world samples were reserved exclusively for testing—enabling evaluation of out-of-distribution (OOD) generalization in authentic CAD scenarios.
+We design four distinct prompt configurations to probe their orthographic projection reasoning capabilities: 1) Test image‐Only: Models receive only the target rasterized drawings; 2) Test image + Reference Image: In addition to the target drawings, models are given a reference image—rendered on the same geometry layer—with all dimension names to supply semantic alignment; 3) Test image + Answered Pair: Inputs consist of the target drawings plus a correctly paired raster image and its parameter vector, providing an exemplar mapping between visual features and quantitative parameters; 4) Test image + Attribute Explanation: Models are supplied with the target drawings alongside a detailed textual interpretation authored by professional engineers.
+Given that orthographic projection reasoning involves complex multi-step inference and current pretrained VLMs have limited exposure to CAD drawings, we incorporated explicit reasoning guidance into each prompt format to encourage coherent reasoning trajectories.
+For each of the four prompt configurations, we outlined step-by-step instructions for task completion—structured to guide the model through the reasoning process systematically.
 The complete set of instructions for each format is provided in the supplementary materials.
@@
 Qwen2.5-VL [42] 24.54 30.76 30.47 25.86 24.54 32.78 33.64 38.88
+Gemini2.5 Pro [43] 24.23 30.09 29.39 25.55 25.68 30.61 32.30 35.88
+Claude4 Vision [44] 23.47 29.67 30.04 25.58 24.94 33.00 34.13 36.32
+GPT-4o  [45] 26.06 31.56 32.28 27.52 27.79 34.17 39.71 38.56
 Ours 80.86 82.99 83.24 82.67 81.35 83.11 82.87 84.03
-5.1 Performance on in-domain test set.
-The experimental results, as shown in Tab.
-3, demonstrate that our method significantly outperforms existing VLMs on orthographic projection reasoning tasks across all prompt formats.
-Specifically, in the Without Reasoning Guidance setting, the "+Answered Pair" prompt achieves the highest accuracy of 83.24%.
-In the Reasoning Guidance setting, our model achieves an accuracy of 84.03% in the "+Attribute Explanation" prompt.
-Furthermore, our method exhibits strong performance even in the "Test Image Only" condition, where we achieve 80.86%.
-This improvement highlights the effectiveness of our approach.
+5.1 Performance on in-domain test set
+The experimental results (Tab.
+3) demonstrate that our method significantly outperforms existing vision–language models (VLMs) on orthographic projection reasoning tasks across all prompt formats.
+Specifically, in the "Without Reasoning Guidance" setting, the "+Answered Pair" prompt achieves the highest accuracy of 83.24%.
+In the "With Reasoning Guidance" setting, our model reaches 84.03% accuracy with the "+Attribute Explanation" prompt.
+Notably, our method also delivers strong performance under the "Test Image Only" condition, achieving 80.86% accuracy—highlighting the effectiveness of our proposed approach.
 It is worth noting that the performance improvement from "without reasoning guidance" to "with reasoning guidance" is not as significant.
@@
 Further analysis regarding the three findings are presented in the supplementary materials.
-5.2 Performance on real-world test set.
-We utilize all 3,000 real-world samples to evaluate out-of-distribution (OOD) generalization in real-world scenarios.
-The baseline model, Qwen2.5-VL, achieves an accuracy of 13.47%, which is lower than its performance on synthetic data (24.54%).
-This performance drop is primarily due to the more complex structure and occlusion relationships present in the real-world data, as discussed in Sec.
-3.1, which increases the difficulty of the task.
-We also compare the results with a model trained using a SFT strategy, which achieve an accuracy of 36.15%.
-This improvement can be attributed to significant gains in Primitive Counting tasks.
-However, the model still struggle with Composite Parameter Computation, highlighting the limitations of traditional SFT.
-Our method, using CReFT, achieves an accuracy of 46.67%, demonstrating that our method can effectively address OOD generalization and lead to significant improvements on more complex real-world data.
+5.2 Performance on real-world test set
+We evaluate out-of-distribution (OOD) generalization using all 3,000 real-world samples, which reflect authentic CAD scenarios.
+The baseline model (Qwen2.5-VL) achieves 13.47% accuracy—lower than its 24.54% performance on synthetic data.
+This performance degradation is primarily attributed to the more complex structures and occlusion relationships in real-world data (Sec.
+3.1), which elevates task difficulty.
+We also compare with a model trained via standard supervised fine-tuning (SFT), which achieves 36.15% accuracy.
+This improvement stems from substantial gains in Primitive Counting tasks.
+However, the model still struggles with Composite Parameter Computation—underscoring the limitations of conventional SFT.
+Our CReFT-based method achieves 46.67% accuracy, demonstrating its ability to effectively mitigate OOD challenges and deliver substantial performance gains on complex real-world data.
 Table 2: Ablation study results.
@@
 2 reveal a significant improvement in performance when incorporating chain-of-thought (CoT) reasoning.
-Specifically, the accuracy for individual Task 3 increase from 46.16% to 74.15% after adding CoT , demonstrating a clear performance boost.
+Specifically, the accuracy for individual Task 3 increases from 46.16% to 74.15% after adding CoT , demonstrating a clear performance boost.
 When combining multiple training tasks, such as Task 1 + Task 3 and Task 2 + Task 3, the improvements are even more pronounced, with accuracies reaching 77.90% and 67.75%, respectively.
@@
 In future work, we plan to explore techniques to enhance the model’s robustness and adaptability.
+ACKNOWLEDGMENT This work is supported in part by NSFC Project (No. 62176061) and Joint Laboratory of Intelligent Construction Engineering Technology for Operating Railway Lines.
 Appendix A The complete reasoning guidance for each input format
@@
 Overall, orthographic projection reasoning entails not only reading textual dimensions and visual features, but also matching annotations to geometry primitives, counting structural elements, and computing composite parameters.
-Consequently, none of the off-the-shelf models fully masters this composite task.
+Consequently, none of the off‑the‑shelf models fully masters this composite task.
 Qwen2.5-VL [46] achieves the highest accuracy—38.88% under reasoning guidance with the Test Image + Attribute Explanation prompt.
 This improvement stems primarily from its superior ability to parse geometry layers and read dimension labels.
-Crucially, its performance on reasoning-intensive parameters remains low.
-These findings underscore that orthographic projection CAD cannot be solved by prompting alone and requires dedicated fine-tuning strategy.
+Crucially, its performance on reasoning‑intensive parameters remains low.
+These findings underscore that orthographic projection CAD cannot be solved by prompting alone and requires dedicated fine‑tuning strategy.
 2) Prompt Format–Dependent Performance Gains.
@@
 3)Introducing reasoning guidance yields consistent gains across all seven VLMs.
-Introducing reasoning guidance demonstrates that step-wise reasoning guidance helps models better decompose the multi-step orthographic projection tasks.
+Introducing reasoning guidance demonstrates that step‐wise reasoning guidance helps models better decompose the multi‐step orthographic projection tasks.
 In the absence of reasoning guidance, visual exemplar prompts (+Reference Image and +Answered Pair) deliver the largest relative improvements.
@@
 It stems from reasoning guidance is textual form, which synergizes most effectively with textual inputs.
-Hence, models with stronger text encoders (e.g., DeepSeek-V1) exhibit disproportionately larger boosts.
+Hence, models with stronger text encoders (e.g., DeepSeek‑V1) exhibit disproportionately larger boosts.
 These results underscore the necessity of integrating structured reasoning guidance to advance orthographic projection CAD reasoning.